OpenStudy (anonymous):

An aqueous solution of NaCl is 350ppm (parts per million). What is the mass of NaCl in 1.00L of the solution? Assume the density of the solution is 1.00g/ml

OpenStudy (anonymous):

Density = m/v 1.00g/ml = m/ 1000ml m=1000g I don't know what to do with this info..